OpenStudy (anonymous):

4t(t-5)-2t(3t+2)+108=3t(2t-5)-t(8t-3)

OpenStudy (anonymous):

\[4t(t-5)-2t(3t+2)+108~~=~~3t(2t-5)-t(8t-3)\]\[~~4t^2-20t-6t^2-4t+108~~=~~6t^2-15t-8t^2+3t\]\[~~~~~~~~~~~~~~-2t^2-24t+108~~=~~-2t^2-12t\]\[~~~~~~~~~~~~~~~~~~~~~~~-12t+108~~=~~0\]\[~~~~~~~~~~~~~~~~~~~~~~~~-12(t-9)~~=~~0\]\[~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~12~~\neq~~0\]\[~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~t-9~~=~~0\]\[\large~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~t=9\]